You are working to earn enough money for a vacation which costs 1400.0 dollars. Your take-home pay is $16.50/hour, but you only work 4-hour shifts each day. How many days will it take for you to earn enough money to pay for your vacation if all the money you earn goes towards your vacation?


1
Expert's answer
2021-01-21T03:51:16-0500

cost of vacation = 1400 dollars

▪home pay = 16.50 dollars per hour

▪total working hours each day = 4

therefore ,

▪total earning each day

16.50 × 4 = 66

▪assume that ,total days required to earn cost of vacation are as x.

▪ X × 66 = 1400

▪ X = "\frac{1400}{66}"

X= 21.2121 days

To earn cost vacation i.e-1400 dollars. Required 21.2121 days